Och, Inc., is considering a project that will result in initial aftertax cash savings of $1.86 million at the end of the first year, and these savings will grow at a rate of 2 percent per year indefinitely. The company has a target debt-equity ratio of .8, a cost of equity of 12.6 percent, and an aftertax cost of debt of 5.4 percent. The cost-saving proposal is somewhat riskier than the usual projects the firm undertakes; management uses the subjective approach and applies an adjustment factor of +3 percent to the cost of capital for such risky projects. What is the maximum initial cost the company would be willing to pay for the project? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89)
Answer #1
D/A = D/(E+D)
D/A = 0.8/(1+0.8)
=0.4444
Weight of equity = 1-D/A
Weight of equity = 1-0.4444
W(E)=0.5556
Weight of debt = D/A
Weight of debt = 0.4444
W(D)=0.4444
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.4*(1-0)
= 5.4
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.4*0.4444+12.6*0.5556
WACC =9.4%

WACC for project = 9.4+3=12.4%

Max initial cost = FCF in 1 year/(WACC - growth rate)
= 1860000/ (0.124 - 0.02)
= 17884615.38